Grade Select GradePhysical Chemistry

a gas is heated through 10 in a closed vessel , so the pressure increases by 0.4%. the initial temp of the gas was

Mrinal Sharma
11 Years agoGrade Select Grade

1 Answer

Pankaj
11 Years ago
Using Ideal Gas Equation
PV = nRT

Here
V, n and R are constants
thus, P and T are proportinal to each other

P2 = T2/T1 * Pold = 1.004 * P1

thus, 1.004 = T2/T1

we also know Tnew = T1+ 1

(T1+1)/(T2) = 1.004
T1 +1 = 1.004 T1
.0096 T1 = 1
T1 = 1/.0096 = 104 K = 104+273 = 381 C